OpenStudy (anonymous):

Let S be a set, and * be a binary operation on S that satisfies two laws: x*x = x for all x in S, and (x*y)*z = (y*z)*x for all x, y, z in S. Prove that * is commutative and associative.

OpenStudy (anonymous):

I think i have an idea of how to prove it, I just want to see what other people have to say. Just for clarification purposes, We want to prove that: x*y = y*x and (x*y)*z = x*(y*z)

OpenStudy (anonymous):

Groupy stuff...

OpenStudy (anonymous):

yeah >.< i thought i had an idea, but as I was typing it, i saw a hole in the logic and it fell apart =/

OpenStudy (anonymous):

If it was a*a it would be GA.

OpenStudy (anonymous):

x*x I mean...

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

if i end up with a statement like: (x*y)*x = (y*x)*x does that imply that x*y = y*x ?

OpenStudy (anonymous):

Yes because of x*x

OpenStudy (anonymous):

the xyx stuff IS associative, isn't it?

OpenStudy (anonymous):

it has to be shown. just from that second law its not clear. My plan is to prove commutative with just those two laws, then prove associative with commutative + two laws.

OpenStudy (anonymous):

Whatdy'u call it, bilinear wotsit..

OpenStudy (anonymous):

if (x*y)*x = (y*x)*x proves x*y = y*x, then thats commutative down. Then associative shouldnt be too bad...i think lolol

OpenStudy (anonymous):

it just seems a little iffy to me >.<

OpenStudy (anonymous):

That must be right if x*x = x for all x.

OpenStudy (anonymous):

How so? im a little confused.

OpenStudy (anonymous):

Oh, I see, moment, anticommutative is still possible....

OpenStudy (anonymous):

I guess i should state how i got that statement. I started with x*y, and did: x*y = (x*x)*y Then by law 2 we can change the order: (x*x)*y = (x*y)*x Using law 2 once more: (x*y)*x = (y*x)*x

OpenStudy (anonymous):

From here I want to say something like, "because binary operations are one-to-one, and the object x is on the right side of both statements, this implies that x*y = y*x, yada yada yada".

OpenStudy (anonymous):

I think this is correct. "If * is a binary operation, and a*b = a*c, then b = c."

OpenStudy (anonymous):

Transitivity.

OpenStudy (anonymous):

i need to verify that though >.< ima search the web lol

OpenStudy (anonymous):

I really much prefer to use the algebras than prove that the algebra exists, lol

OpenStudy (anonymous):

me too >.< this is a for a problem solving class. We just pick random problems to solve. This one looked do-able so i picked it.

OpenStudy (anonymous):

I'm just thinking because GA is noncommutative algebra but associative and I am trying to see what exactly it is in yours makes it commutative and it can only be the x*x = x

OpenStudy (anonymous):

Whereas in GA it's x^2

OpenStudy (anonymous):

That allows u to pull that substitution trick...

OpenStudy (anonymous):

right right. I think i have enough info to write out the proof now. thanks :)

OpenStudy (anonymous):

Pick something less, um... well u know, next ime, lol:-)

OpenStudy (anonymous):

Alright, i came up with better reasoning. This looks so much better lol. Start with x*y. Using property 1 we get: x*y = (x*y)*(x*y) Then by property 2 we have: (x*y)*(x*y) = [(x*y)*x]*y Using property 2 on the inside argument backwards we obtain: [(x*y)*x]*y = [(x*x)*y]*y = [x*y]*y Then one more property 2 gives: [x*y]*y = (y*y)*x = y*x Thus x*y = y*x, and * is commutative. For Associative, we start with property 2: (x*y)*z = (y*z)*x Then using the commutative property we just proved, we obtain: (y*z)*x = x*(y*z) So (x*y)*z = x*(y*z), and * is associative.